A force of 500n acts over an area of 10m2 .use the formula to calculate the pressure
1
Expert's answer
2020-04-28T09:41:27-0400

The pressure by definition


"P=\frac{F}{A}=\frac{500\:\rm N}{10\:\rm m^2}=50\:\rm Pa."